# Mixture of Experts-augmented Deep Unfolding for Activity Detection in IRS-aided Systems

### v1

@@ I Introduction
Massive machine-type communications (mMTC) have been expected to play a vital role to empower the sixth-generation (6G) vision of future ubiquitous connectivity. To meet the low latency requirement in mMTC, grant-free random access is recognized as a promising solution. However, grant-free random access requires the base station (BS) to perform activity detection [1, 2, 3]. Due to the large number of potential Internet of Things (IoT) devices and nonorthogonal signature sequences, activity detection is a challenging task.
@@
On the other hand, deep learning-based algorithms have become popular in communication research due to their ability to overcome modeling inaccuracy [11, 12]. However, black box deep learning designs (e.g., multi-layer perceptrons (MLP) or convolutional neural networks) do not incorporate the domain knowledge of communication systems. This usually results in unsatisfactory performance or the neural network not being able to converge during training. To overcome such drawbacks, model-driven deep unfolding has emerged as a viable alternative. By regarding each iteration of an optimization algorithm as one layer of the neural networks, the model-driven deep unfolding algorithm embeds the domain knowledge into neural network designs while exploiting the data to determine the behavior of the model.
@@ II-A System Model
Denote the device activity indicator of device $k$ as $b_{k}=1$ if it is active and 0 otherwise. Moreover, a unique signature sequence ${\mathbf{s}}_{k}=\left[s_{k,1},s_{k,2},\ldots,s_{k,L}\right]^{T}\in\mathbb{C}^{L}$ is assigned to each device $k$. At the start of each block, all the active devices send their signature sequences and the BS conducts activity detection based on the received signals. Assuming that the transmissions from different devices are synchronous, the received signal is derived as
@@
II-B Problem Formulation
The activity detection problem is mathematically equivalent to detecting whether $a_{k}\triangleq b_{k}\sqrt{p_{k}\beta_{k}}$ is positive or zero. To this end, we treat the device activities $\{a_{k}\}_{k\in\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}}$ as a set of unknown but deterministic parameters, and model the received signal $\mathbf{Y}$ as a random variable. With $\mathbf{y}_{m}$ denoting the $m$-th column of $\mathbf{Y}$, the covariance matrix of $\mathbf{y}_{m}$ is given by the following proposition.
Proposition 1.
[…]
@@ II-B Problem Formulation
where
$\bar{\mathbf{y}}_{m}&=\sum_{k\in\mathcal{K}_{1}}{a_{k}}\mathbf{s}_{k}(\sqrt{\frac{\kappa_{\text{U}}}{1+\kappa_{\text{U}}}}\mathbf{h}_{k}^{\text{LoS}})^{H}\mathbf{\Theta}\sqrt{\frac{\kappa_{\text{B}}}{1+\kappa_{\text{B}}}}\mathbf{g}_{m}^{\text{LoS}}\\ &+\sum_{k\in\mathcal{K}_{2}}{a_{k}}\mathbf{s}_{k}\sqrt{\frac{\kappa_{\text{U}}}{1+\kappa_{\text{U}}}}\mathbf{f}_{k}^{\text{LoS}}(m)^{*}\end{aligned}$ (6)
and
$\boldsymbol{\Sigma}_{m}=\sum_{k\in\mathcal{K}_{1}}{a_{k}}\Xi_{m,k}\mathbf{s}_{k}\mathbf{s}_{k}^{H}&+\sum_{k\in\mathcal{K}_{2}}\frac{a_{k}\mathbf{s}_{k}\mathbf{s}_{k}^{H}}{1+\kappa_{\text{U}}}\\ &+\sum_{k\in\mathcal{K}_{3}}{a_{k}}\mathbf{s}_{k}\mathbf{s}_{k}^{H}+\sigma_{w}^{2}\mathbf{I}_{M},\end{aligned}$ (7)
with $\mathbf{f}_{k}^{\text{LoS}}(m)$ denoting the $m$-th element of $\mathbf{f}_{k}^{\text{LoS}}$ and $\Xi_{m,k}=\frac{N+\kappa_{\text{B}}\|{\mathbf{g}}_{m}^{\text{LoS}}\|_{2}^{2}+\kappa_{\text{U}}\|{\mathbf{h}}_{k}^{\text{LoS}}\|_{2}^{2}}{(1+\kappa_{\text{B}})(1+\kappa_{\text{U}})}$.
Proof.
Please see Appendix A. ∎
@@ III-A Projected Gradient Dscent Approach
$d_{k}^{(i-1)}=\sum_{m=1}^{M}d_{k,m}^{(i-1)}$ (10)
where $d_{k,m}^{(i-1)}$ is given on the top of this page. Algorithm 1 summarizes the PGD approach.
@@
1:  Input The received signal ${\mathbf{Y}}$, the signature matrix ${\mathbf{S}}$, phase shift $\mathbf{\Theta}$, and the noise power $\sigma_{{w}}^{2}$.
2:  Initialize $\left\{\hat{a}_{k}^{(0)}=0\right\}_{k\in\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}}$.
3:  repeat ($i=1,2,...$)
4:      Calculate $\boldsymbol{\Sigma}_{m}^{(i-1)}$ by (7).
5:      Calculate the gradient $d_{k}^{(i-1)}$ at $\hat{a}_{k}^{(i-1)}$ by performing $d_{k}^{(i-1)}=\sum_{m=1}^{M}d_{k,m}^{(i-1)}$.
6:      Calculate $\hat{a}_{k}^{(i)}=\max\left\{\hat{a}_{k}^{(i-1)}-\eta^{(i)}d_{k}^{(i-1)},0\right\},~{}\forall k\in\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}$.
7:  until Convergence
8:  Output $\left\{\hat{a}^{(I)}_{k}\right\}_{k\in\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}}$ where $I$ is the last iteration index.
Algorithm 1 PGD
@@ III-B The Deep Unfolded Network
Applying deep unfolding to the PGD approach, each iteration of Algorithm 1 is regarded as one layer of the neural network. The key in constructing a deep unfolded network is to introduce trainable parameters in each layer of the neural network. In Algorithm 1, since the choice of $\eta^{(i)}$ influences the convergence rate and detection performance, we regard it as a trainable parameter. Furthermore, to avoid the matrix inverse in line 5 of Algorithm 1 and increase the learning capability, we design a linear layer $\mathbf{A}^{(i)}\boldsymbol{\Sigma}_{m}^{(i-1)}+\mathbf{B}^{(i)}$ with trainable parameters $\mathbf{A}^{(i)}$ and $\mathbf{B}^{(i)}$ to approximate $\left(\boldsymbol{\Sigma}_{m}^{(i-1)}\right)^{-1}$.
The unrolled network consists of $I$ cascaded layers with trainable parameters $\left\{\mathbf{A}^{(i)},\mathbf{B}^{(i)},\eta^{(i)}\right\}_{i=1}^{I}$ as shown in Fig. 2. Specifically, for the $i$-th layer, the input is $\left(\mathbf{Y},\mathbf{S},\left\{\hat{a}_{k}^{(i-1)}\right\}_{k\in\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}}\right)$ and the output is $\left\{\hat{a}_{k}^{(i)}\right\}_{k\in\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}}$.
@@ III-C Mixture of Experts Design
Both the PGD and its deep unfolded version require the computation of $\boldsymbol{\Sigma}_{m}^{(i-1)}$ in (7) with the current activity status $\hat{a}_{k}^{(i-1)}$. However, calculating $\boldsymbol{\Sigma}_{m}^{(i-1)}$ requires the precise knowledge of which devices belong to which group: $\mathcal{K}_{1}$, $\mathcal{K}_{2}$, or $\mathcal{K}_{3}$. This information is difficult to acquire in practice. One naive way to handle this is to assume all the devices belong to only one of the groups. This corresponds to the covariance matrix being
@@
However, how do we know which covariance matrix we should use in practice? To address this challenge, we propose a MoE design. In particular, one of the major challenges in constructing a MoE architecture lies in designing the gate network to process input information and assigning the most appropriate expert for subsequent computation. To this end, we propose a MLP as the gate network. Upon feeding the received signal matrix ${\mathbf{Y}}$ into the gate network, it outputs the proportions of three types of devices. By detecting the device type with the highest proportion, we select the corresponding expert to compute the covariance matrix. The specific structure of the MoE system and its integration with the unfolded network are illustrated in Fig. 2. By introducing this MoE-module, it is not necessary to acquire the channel type information between each device and the BS anymore.
@@ IV Simulation Results
In this section, we compare the performance of the proposed MoE-augmented deep unfolding approach, traditional coordinate descent (CD) algorithm (assume all channels are zero-mean Gaussian), the derived PGD algorithm and a transformer-based detector [15]. After obtaining the estimate $\hat{a}_{k}$ for each device, the estimate indicator $\hat{b}_{k}$ is recovered from
$\hat{b}_{k}=\begin{cases}1,&\text{if}\quad\hat{a}_{k}\geq a^{th},\\ 0,&\text{else},\end{cases}$ (11)
where $a^{th}$ is a threshold that controls the trade-off between the probabilities of false alarm (PF) and missed detection (PM), which are defined as [6]:
$\mathrm{PM}=1-\frac{\sum_{k=1}^{K}{b}_{k}\hat{b}_{k}}{\sum_{k=1}^{K}{b}_{k}},~{}~{}\mathrm{PF}=\frac{\sum_{k=1}^{K}\hat{b}_{k}\left(1-{b}_{k}\right)}{\sum_{k=1}^{K}\left(1-{b}_{k}\right)}.$
IV-A Simulation Setting and Training Procedure
@@
Under a three dimensional Cartesian coordinate system, we set the locations of the BS and the IRS at $(0,0,10)$ and $(5,50,10)$ in meter, respectively. In addition, devices from $\mathcal{K}_{1}$ are randomly and uniformly located in a circular area of radius $40$ m around the center $(200,0,0)$. And the devices in $\mathcal{K}_{2}$ and $\mathcal{K}_{3}$ are located in the left half circular area with radius $40$m around the center $(0,120,0)$. The large-scale coefficient of each device $\beta_{k}$ is generated according to $\beta_{k}=-60-22\log_{10}(d_{k}d_{0})$ in dB, where $d_{k}$ denotes the distance between device $k$ and the IRS, and $d_{0}$ denotes the distance between the IRS and the BS. The transmit power of each device is set as $p_{k}=23$ dBm. The BS is equiped with $32$ antennas and the IRS is with $40$ reflecting elements. The number of potential devices is $K=100$, with the activity probability being 0.2. The length of signature sequence $s_{k}$ is set at $L=20$.
@@
To train an MLP-based gate network, the received signal matrix $\mathbf{Y}\in\mathbb{C}^{L\times M}$ is first preprocessed by concatenating real and imaginary components into a real-valued vector $\mathbf{y}_{\text{real}}=[\text{Re}(\text{vec}(\mathbf{Y}));\text{Im}(\text{vec}(\mathbf{Y}))]\in\mathbb{R}^{2LM}$. The MLP architecture consists of three fully-connected layers: an input layer ($2LM\to 512$) with ReLU activation, a hidden layer ($512\to 128$) with tanh activation, and an output layer ($128\to 3$) with softmax normalization to produce $\rho_{1}$, $\rho_{2}$, $\rho_{3}$ as the estimate of user group proportions. The training dataset composes of $1.28\times 10^{6}$ training samples and the gate network is trained using KL divergence loss:
$\mathcal{L}=\sum_{i=1}^{3}\rho_{i}^{\text{true}}\log\left(\frac{\rho_{i}^{\text{true}}}{\rho_{i}}\right),$ (12)
where
$\rho_{i}^{\text{true}}=\frac{|\mathcal{K}_{i}|}{\sum_{i=1}^{3}|\mathcal{K}_{i}|},\quad i=1,2,3.$ (13)
For the deep unfolding network, we use incremental training [16], and it is found that the network with more than 4 layers only leads to a small performance improvement. Therefore, we choose $I=4$.
IV-B Simulation Results
First, we present the PM-PF curves of various methods under the setting of: $\kappa_{\text{U}}$ = $\kappa_{\text{R}}$ = $\kappa_{\text{B}}$ = 10 dB, $\sigma_{w}^{2}$ = -95 dBm. Among 100 devices, 40% of them are from $\mathcal{K}_{1}$ and the proportions of devices in $\mathcal{K}_{2}$ and $\mathcal{K}_{3}$ are both 30%. In Fig. 3, it can be observed that the proposed deep unfolding design significantly outperforms both the optimization-based PGD and CD methods as the deep unfolded network could compensate the inaccurate modeling. Furthermore, the deep unfolding method is superior than the transformer based detector, since deep unfolding utilizes the mathematical principles as implicit training guidelines, making it easier to achieve better detection performance. While the original deep unfolding method still needs to obtain the specific channel type information between devices and the BS, the MoE-aided deep unfolding network, which does not need the knowledge of which device belongs to which group, only suffers from a small degradation compared to the deep unfolded network with perfect information. In contrast, if no such information is available for the PGD algorithm, it has a significant degradation compared to the PGD with perfect information.
@@ IV-B Simulation Results
Next, Fig. 4 illustrates the flexibility and generalization ability of the proposed MoE-augmented deep unfolding approach. We compared the error rates of various methods with the proportion of devices in $\mathcal{K}_{1}$ varies from 0% to 80%. The number of remaining devices is equally divided between $\mathcal{K}_{2}$ and $\mathcal{K}_{3}$. We can observe that the deep unfolding network that uses a single fixed expert only performs well if the network setting matches the expert’s prior knowledge. However, with the MoE system enabled, the gate network can learn which type of device has the highest proportion and select the best expert accordingly. This property leads to the MoE-based deep unfolding network performing close to the deep unfolding network with perfect devices’ grouping information.

### v2

@@ I Introduction
Massive machine-type communications (mMTC) have been expected to play a vital role to empower the sixth-generation (6G) vision of future ubiquitous connectivity [1]. To meet the low latency requirement in mMTC, grant-free random access is recognized as a promising solution. However, grant-free random access requires the base station (BS) to perform activity detection [2, 3, 4]. Due to the large number of potential Internet of Things (IoT) devices and nonorthogonal signature sequences, activity detection is a challenging task.
@@
On the other hand, deep learning-based algorithms have become popular in communication research due to their ability to overcome modeling inaccuracy [12, 13]. However, black box deep learning designs (e.g., multi-layer perceptrons (MLP) or convolutional neural networks) do not incorporate the domain knowledge of communication systems. This usually results in unsatisfactory performance or the neural network not being able to converge during training. To overcome such drawbacks, model-driven deep unfolding has emerged as a viable alternative. By regarding each iteration of an optimization algorithm as one layer of the neural networks, the model-driven deep unfolding algorithm embeds the domain knowledge into neural network designs while leveraging on the data to determine the behavior of the model.
@@ II System Model and Problem Formulation
Denote the device activity indicator of device $k$ as $b_{k}=1$ if it is active and 0 otherwise. Moreover, a unique signature sequence ${\mathbf{s}}_{k}=\left[s_{k,1},s_{k,2},\ldots,s_{k,L}\right]^{T}\in\mathbb{C}^{L}$ is assigned to each device $k$. At the start of each block, all the active devices send their signature sequences and the BS conducts activity detection based on the received signals. Assuming that the transmissions from different devices are synchronous, the received signal is
@@
The activity detection problem is mathematically equivalent to detecting whether $a_{k}\triangleq b_{k}\sqrt{p_{k}\beta_{k}}$ is positive or zero. To this end, we treat the device activities $\{a_{k}\}_{k\in\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}}$ as a set of unknown but deterministic parameters, and model the received signal $\mathbf{Y}$ as a random variable. In particular, denoting the the non-line-of-sight parts of (2) and (3) as $\mathbf{h}_{k}^{\text{NLoS}}\sim\mathcal{CN}\left(\mathbf{0},\mathbf{I}_{N}\right)$ and $\mathbf{g}_{m}^{\text{NLoS}}\sim\mathcal{CN}\left(\mathbf{0},\mathbf{I}_{N}\right)$. Based on [16, Proposition 1], we have $(\mathbf{h}_{k}^{\text{NLoS}})^{H}\mathbf{\Theta}\mathbf{g}_{m}^{\text{NLoS}}\sim\mathcal{CN}\left(\mathbf{0},N\mathbf{I}_{M}\right)$ when $N$ is sufficiently large. Therefore, $\mathbf{Y}$ is a complex Gaussian distributed matrix. The mean of the $m$-th column of $\mathbf{Y}$ (denoted as $\mathbf{y}_{m}$) is then given by
$\bar{\mathbf{y}}_{m}\triangleq\mathbb{E}\left[\mathbf{y}_{m}\right]$ $=\sum_{k\in\mathcal{K}_{1}}{a_{k}}\mathbf{s}_{k}(\sqrt{\frac{\kappa_{\text{R}}}{1+\kappa_{\text{R}}}}\mathbf{h}_{k}^{\text{LoS}})^{H}\mathbf{\Theta}\sqrt{\frac{\kappa_{\text{B}}}{1+\kappa_{\text{B}}}}\mathbf{g}_{m}^{\text{LoS}}$
$+\sum_{k\in\mathcal{K}_{2}}{a_{k}}\mathbf{s}_{k}\sqrt{\frac{\kappa_{\text{U}}}{1+\kappa_{\text{U}}}}\mathbf{f}_{k}^{\text{LoS}}(m)^{*},$ (5)
where $\mathbf{f}_{k}^{\text{LoS}}(m)$ denotes the $m$-th element of $\mathbf{f}_{k}^{\text{LoS}}$. Furthermore, the covariance of $\mathbf{y}_{m}$ can be computed as
$\boldsymbol{\Sigma}_{m}\triangleq\mathbb{E}\left[(\mathbf{y}_{m}-\mathbb{E}\left[\mathbf{y}_{m}\right])(\mathbf{y}_{m}-\mathbb{E}\left[\mathbf{y}_{m}\right])^{H}\right]$
$=\!\sum_{k\in\mathcal{K}_{1}}{a_{k}}\Xi_{m,k}\mathbf{s}_{k}\mathbf{s}_{k}^{H}\!+\!\!\!\sum_{k\in\mathcal{K}_{2}}\frac{a_{k}\mathbf{s}_{k}\mathbf{s}_{k}^{H}}{1+\kappa_{\text{U}}}+\!\!\!\sum_{k\in\mathcal{K}_{3}}{a_{k}}\mathbf{s}_{k}\mathbf{s}_{k}^{H}+\sigma_{w}^{2}\mathbf{I}_{M},$ (6)
where $\Xi_{m,k}=\frac{N+\kappa_{\text{B}}\|{\mathbf{g}}_{m}^{\text{LoS}}\|^{2}+\kappa_{\text{U}}\|{\mathbf{h}}_{k}^{\text{LoS}}\|^{2}}{(1+\kappa_{\text{B}})(1+\kappa_{\text{U}})}$. Once obtaining the mean and covariance of $\mathbf{y}_{m}$, the likelihood function of $\left\{a_{k}\right\}_{k\in\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}}$ is given by
@@
where $L$ denotes the length of signature sequence.
@@ III-A Projected Gradient Dscent Approach
$d_{k}^{(i-1)}=\sum\nolimits^{M}_{m=1}d_{k,m}^{(i-1)}$ (10)
where $d_{k,m}^{(i-1)}$ is given on the top of this page.
@@ III-B The Deep Unfolded Network
Applying deep unfolding to the PGD approach, each iteration of (9) and (10) is regarded as one layer of the neural network. The key in constructing a deep unfolded network is to introduce trainable parameters in each layer of the neural network. In the PGD algorithm, since the choice of $\eta^{(i)}$ influences the convergence rate and detection performance, we regard it as a trainable parameter. Furthermore, to avoid the matrix inverse in the computation of $d_{k}^{(i-1)}$ and increase the learning capability, we design a linear layer $\mathbf{A}^{(i)}\boldsymbol{\Sigma}_{m}^{(i-1)}+\mathbf{B}^{(i)}$ with trainable parameters $\mathbf{A}^{(i)}$ and $\mathbf{B}^{(i)}$ to approximate $(\boldsymbol{\Sigma}_{m}^{(i-1)})^{-1}$.
The unfolded network consists of $I$ cascaded layers with trainable parameters $\left\{\mathbf{A}^{(i)},\mathbf{B}^{(i)},\eta^{(i)}\right\}_{i=1}^{I}$ as shown in Fig. 2. Specifically, for the $i$-th layer, the input is $(\mathbf{Y},\mathbf{S},\{\hat{a}_{k}^{(i-1)}\}_{k\in\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}})$ and the output is $\{\hat{a}_{k}^{(i)}\}_{k\in\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}}$.
@@ III-C Mixture of Experts Design
Both the PGD and its deep unfolded version require the computation of $\boldsymbol{\Sigma}_{m}^{(i-1)}$ in (II) with the current activity status $\hat{a}_{k}^{(i-1)}$. However, calculating $\boldsymbol{\Sigma}_{m}^{(i-1)}$ requires the precise knowledge of which devices belong to which group: $\mathcal{K}_{1}$, $\mathcal{K}_{2}$, or $\mathcal{K}_{3}$. This information is difficult to acquire in practice. One naive way to handle this is to assume all the devices belong to only one of the groups. This corresponds to the covariance matrix being
@@
However, how do we know which covariance matrix we should use in practice? To address this challenge, we propose a MoE design. However, one of the major challenges in constructing a MoE architecture lies in designing the gate network to process input information and assigning the most appropriate expert for subsequent computation. To this end, we propose a MLP as the gating network. Upon feeding the received signal matrix ${\mathbf{Y}}$ into the gating network, it outputs the proportions of three types of devices. By detecting the device type with the highest proportion, we select the corresponding expert to compute the covariance matrix. The specific structure of the MoE system and its integration with the unfolded network are illustrated in Fig. 2. By introducing this MoE-module, it is not necessary to acquire the channel type information between each device and the BS anymore.
The MoE framework aligns with the mixture-of-experts theory [17, 18], where the gating network learns to partition the input space into regions dominated by specific experts. By minimizing the KL divergence loss:
$\mathcal{L}=\sum\nolimits_{i=1}^{3}\rho_{i}^{\text{true}}\log\left({\rho_{i}^{\text{true}}}/{\rho_{i}}\right),$ (11)
the gating network asymptotically approximates the true device group proportions,
$\rho_{i}^{\text{true}}={|\mathcal{K}_{i}|}/{\sum\nolimits_{i=1}^{3}|\mathcal{K}_{i}|},\quad i=1,2,3.$ (12)
This ensures that the selected expert matches the dominant channel type, thereby adapting to mixed fading conditions. As analyzed in [17], this nonlinear top-1 expert selection MoE performs well on classification problems.
The MoE-augmented deep unfolding network is trained end-to-end using incremental training [16]. The gating network’s stability is ensured by the convexity of the KL divergence loss, which guarantees that gradient-based updates converge to a local minimum [18]. The overall network convergence is further guaranteed by the convergence properties of the PGD algorithm [16].
The computational complexity of the deep unfolding network without the MoE module is $\mathcal{O}(I\cdot(M^{3}+LM))$. And with the MoE module, the computational complexity becomes $\mathcal{O}(LM+I\cdot(M^{3}+LM))$. Given that $LM\ll I\cdot M^{3}$ (with $L=20$, $M=32$ and $I=4$ in the simulation), the impact of the MoE module on the total complexity is negligible.
@@ IV Simulation Results
In this section, we compare the performance of the proposed MoE-augmented deep unfolding approach with two deep unfolding baselines [19, 20], traditional coordinate descent (CD) algorithm (assume all channels are zero-mean Gaussian), the derived PGD algorithm and a transformer-based detector [21]. After obtaining the estimate $\hat{a}_{k}$ for each device, the estimated activity $\hat{b}_{k}=1$ if $\hat{a}_{k}\geq a^{th}$ (otherwise, $\hat{b}_{k}=0$), where $a^{th}$ is a threshold that controls the trade-off between the probabilities of false alarm $\mathrm{PF}=\frac{\sum_{k=1}^{K}\hat{b}_{k}\left(1-{b}_{k}\right)}{\sum_{k=1}^{K}\left(1-{b}_{k}\right)}$ and missed detection $\mathrm{PM}=1-\frac{\sum_{k=1}^{K}{b}_{k}\hat{b}_{k}}{\sum_{k=1}^{K}{b}_{k}}$ [7].
@@
Under a three dimensional Cartesian coordinate system, we set the locations of the BS and the IRS at $(0,0,10)$ and $(5,50,10)$ in meter, respectively. In addition, devices from $\mathcal{K}_{1}$ are randomly and uniformly located in a circular area of radius $40$ m around the center $(200,0,0)$. The devices in $\mathcal{K}_{2}$ and $\mathcal{K}_{3}$ are located in the left half circular area with radius $40$m around the center $(0,120,0)$. The large-scale coefficient of each device $\beta_{k}$ is generated according to $\beta_{k}=-60-22\log_{10}(\lambda_{k}\lambda_{0})$ in dB, where $\lambda_{k}$ denotes the distance between device $k$ and the IRS, and $\lambda_{0}$ denotes the distance between the IRS and the BS. The transmit power of each device is set as $p_{k}=23$ dBm. The BS is equiped with $32$ antennas and the IRS is with $40$ reflecting elements. The number of potential devices is $K=100$, with the activity probability being 0.2. The length of signature sequence $\mathbf{s}_{k}$ is set at $L=20$.
@@
To train an MLP-based gate network, the received signal matrix $\mathbf{Y}\in\mathbb{C}^{L\times M}$ is first preprocessed by concatenating real and imaginary components into a real-valued vector $\mathbf{y}_{\text{real}}=[\text{Re}(\text{vec}(\mathbf{Y}));\text{Im}(\text{vec}(\mathbf{Y}))]\in\mathbb{R}^{2LM}$. The MLP for MoE consists of three fully-connected layers: an input layer ($2LM\to 512$) with ReLU activation, a hidden layer ($512\to 128$) with tanh activation, and an output layer ($128\to 3$) with softmax normalization to produce $\rho_{1}$, $\rho_{2}$, $\rho_{3}$ as the estimate of user group proportions. For the deep unfolding network, we use incremental training [16], and it is found that the network with more than 4 layers only leads to a small performance improvement. Therefore, we choose $I=4$.
@@
First, we present the PM-PF curves of various methods under the setting of: $\kappa_{\text{U}}$ = $\kappa_{\text{R}}$ = $\kappa_{\text{B}}$ = 10 dB, $\sigma_{w}^{2}$ = -95 dBm. Among 100 devices, 40% of them are from $\mathcal{K}_{1}$ and the proportions of devices in $\mathcal{K}_{2}$ and $\mathcal{K}_{3}$ are both 30%. In Fig. 3, it can be observed that the proposed deep unfolding design significantly outperforms both the optimization-based PGD and CD methods as the deep unfolded network could compensate the inaccurate modeling. Furthermore, the deep unfolding method is superior than the transformer based detector, since deep unfolding utilizes the mathematical principles as implicit training guidelines, making it easier to achieve better detection performance. While the original deep unfolding method still needs specific channel type information between devices and the BS, the MoE-aided deep unfolding network, which does not need the knowledge of which device belongs to which group, only suffers from a small degradation compared to the deep unfolded network with perfect channel type information. In contrast, if no such information is available for the PGD algorithm, it has a significant degradation compared to the PGD with perfect information. When compared with the two unfolding baselines, the proposed unfolding network performs better as it is designed specifically for mixed channel types scenarios while the baselines are designed for one type of channel only.
Next, Fig. 4 illustrates the flexibility and generalization ability of the proposed MoE-augmented deep unfolding approach. We compare the error rates of various methods with the proportion of devices in $\mathcal{K}_{1}$ varies from 0% to 80%. The remaining devices is equally divided between $\mathcal{K}_{2}$ and $\mathcal{K}_{3}$. We can observe that the deep unfolding network that uses a single fixed expert only performs well if the network setting matches the expert’s prior knowledge. However, with the MoE system enabled, the gate network can learn which type of device has the highest proportion and select the best expert accordingly. This property leads to the MoE-based deep unfolding network performing close to the deep unfolding network with perfect devices’ grouping information. For the two deep unfolding baselines, they perform similar to the proposed unfolded PGD method with the covariance matrix fixed as expert 3. This is not surprising as the two baselines also assume all users experience Rayleigh fading.
Finally, we present the computation times of different deep learning methods. As shown in Fig. 5, the MoE-augmented deep unfolding approach experiences the shortest inference time among three deep learning methods. Comparing to the deep unfolding approach with perfect channel type information, the proposed MoE approach, although involves an additional MLP, exhibits even faster computation time since the covariance computation is simpler by focusing on the dominant channel type in the network. Meanwhile, transformer-based detector experiences the slowest inference time due to its generic structure.
Figure 5: Latency versus percentage of devices in $\mathcal{K}_{1}$